OpenStudy (anonymous):

A 135 foot cable is cut into two pieces, one which is 7 feet longer than the other. What is the length of the longer piece?

OpenStudy (anonymous):

\[x+x+7=135\] \[2x+7=135\] \[2x=128\] \[x=64\]

OpenStudy (anonymous):

shorter one is 64, longer is 71

OpenStudy (anonymous):

how is the longer one 71?

OpenStudy (stacey):

If we let x represent the shorter piece of rope, then the longer piece of rope is x+7.

OpenStudy (anonymous):

because i solved for the short one and it was 64. the long one is 7 feet longer than 64 so it must be 71

OpenStudy (anonymous):

you can check that this is right. 71 is 7 more than 64 and 64+71=135

OpenStudy (stacey):

By adding those two pieces we will have the whole cable or 135 feet. So the starting equation was x + (x+7) = 135.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

so, longer cable is of 21 feet and shorter = 14 feet

OpenStudy (anonymous):

what?